OpenStudy (anonymous):

IDK I dont understand Solve this system of equations by using elimination. 3⋅x - 6⋅y = 6 2⋅x + 6⋅y = -6

OpenStudy (anonymous):

can u help/teach me?

OpenStudy (mertsj):

1. Add the two equations. 2. Solve for x 3. Replace x with the value you just got and solve for y.

OpenStudy (anonymous):

yes how do i add them?

OpenStudy (smored):

3x-6y=6 2x+6y=-6 +_________ 5x+0y=-6

OpenStudy (smored):

oops 5x+0y=0

OpenStudy (smored):

so 5x=0 which means x=0

OpenStudy (anonymous):

im doing a study guide i need it right

OpenStudy (smored):

what do you mean?

OpenStudy (anonymous):

i mean i need to get right awnser, but its a (___,___)

OpenStudy (smored):

(x,y) replace the answers with its corresponding variables

OpenStudy (anonymous):

0,1

OpenStudy (smored):

i'd say y=-1 so since you know x=0, substitute it 3(0)-6y=6 -6y=6 y=-1

OpenStudy (anonymous):

ok thank u

OpenStudy (smored):

no prob